write a quadratic equation when the solutions are 3 and 7
Found 2 solutions by solver91311, benazir.sj@gmail.com:
Answer by solver91311(24713)



If is a solution of the quadratic equation , then is a factor of . You have two solutions, so you can write the two factors of your quadratic polynomial.

John

My calculator said it, I believe it, that settles it


Answer by benazir.sj@gmail.com(109)

To get solutions of 3 and 7, then just use (x-3) and (x-7)
(x-3)(x-7)
=x(x-7)-3(x-7)
=x^2-7x -3x+21
x^2-10x +21
the quadratic equation is
x^2-10x+21=0